# There and Back Again: On the relation between Noise and Image Inversions in Diffusion Models

### v2

@@ 4.2 Diversity in conditional generation
Using source prompts $P_{S}$, we generate images $I_{S}$ from Gaussian noise ${\mathbf{x_{T}}}$ and invert them back into the latents ${\mathbf{\hat{x}_{T}}}$. Next, we regenerate images $\hat{I}_{T}$ from the latents, changing the conditioning to the target prompts $P_{T}$. As the ground truth regenerations $I_{T}$, we generate images from the original noise ${\mathbf{x_{T}}}$ with target prompts $P_{T}$. In Fig. 6, we present the drawback of leveraging latent variables as starting points for the denoising, where the $I_{S}$ images’ structures in the latents limit editing performance at those locations in $\hat{I}_{T}$ images.
Figure 6: Uniform regions in the images, encoded in the latents, reduce diversity in the generations when starting from inverse DDIM latents (third row) with different conditioning prompts. In contrast, using clean noise as the starting point (second row) leads to more diverse outputs.
@@
In Tab. 7, we quantitatively measure this effect. First, we calculate the diversity of target generations ($I_{T}$,$\hat{I}_{T}$) against source images ($I_{S}$). We use DreamSim distance [9], LPIPS [56], SSIM [51], and cosine similarity of DINO features [5] to measure the distance between two sets of generations. The experiment shows that edits resulting from latent encodings $\hat{I}_{T}$ are characterized by higher similarity (SSIM, DINO) and lower diversity (DreamSim, LPIPS) relative to starting images $I_{S}$ than the one resulting from noises $I_{T}$. At the same time, in the bottom rows of Tab. 7, we show that the correlations occurring in the latent encodings induce lower performance in text-alignment to target prompts $P_{T}$, which we measure by calculating cosine similarity between text embeddings and resulting image embeddings, both obtained with the CLIP [38] encoder. For the DiT model, our experiment indicates that generations from latents are as similar to the target prompt from the regeneration as to the source prompt used only during the inversion procedure.
@@
Diversity DreamSim $\uparrow$ $0.71_{\pm 0.12}$ $0.69_{\pm 0.13}$ $0.67_{\pm 0.10}$ $0.64_{\pm 0.11}$
against $I_{S}$ LPIPS $\uparrow$ $0.59_{\pm 0.12}$ $0.57_{\pm 0.12}$ $0.39_{\pm 0.11}$ $0.36_{\pm 0.11}$
SSIM $\downarrow$ $0.23_{\pm 0.13}$ $0.25_{\pm 0.14}$ $0.31_{\pm 0.15}$ $0.35_{\pm 0.15}$
DINO $\downarrow$ $0.17$ $0.20$ $0.33$ $0.38$
Text CLIP-T ($P_{S}$) $\downarrow$ $0.456$ $0.543$ $0.0$ $0.0$
alignment CLIP-T ($P_{T}$) $\uparrow$ $0.689$ $0.568$ $0.0$ $0.0$
@@
Table 7: Diversity in relation to source images $I_{S}$ and prompt alignment for generations from noise ${\mathbf{x_{T}}}$ and latents ${\mathbf{\hat{x}_{T}}}$, conditioned on target prompt. The arrows ($\uparrow$/$\downarrow$) indicate greater generation diversity and higher text alignment to the target prompt.
4.3 Noise-to-sample vs. latent-to-sample mapping
As indicated by previous works [19, 54], diffusion models converge to the same mapping between the random Gaussian noise (${\mathbf{x_{T}}}$) and the generated images (${\mathbf{x_{0}}}$) independently on the random seed, parts of the dataset, or even the model architecture. We further investigate this property from the noise-to-sample and sample-to-latent mapping perspective.
We generate $2$K samples from Gaussian noise, and for a given noise ${\mathbf{x_{T}}}$, we predict the resulting generation ${\mathbf{x_{0}}}$ (and vice versa) by calculating its $l_{2}$-distance to all the images and selecting the closest one. We provide more details on the assignment procedure in Sec. E.1.
In Tab. 8, we show that it is possible to accurately assign the starting noises to the corresponding image generations (${\mathbf{x_{0}}}\to{\mathbf{x_{T}}}$) for pixel diffusion models, while reverse assignment (${\mathbf{x_{T}}}\to{\mathbf{x_{0}}}$) is not possible. In Sec. E.3, we explain that assigning images to the noise is unfeasible because of the singular generations with low pixel variance, which are the closest for most of the noises. This fact, connected with the findings on reduced diversity of latents, suggests that latents, unlike noise, cannot be accurately assigned to samples, as evidenced by Tab. 8. It shows that latents significantly differ from noise in their mapping to images.
Model Noise-Image Assignment Latent-Image Assignment
${\mathbf{x_{0}}}\rightarrow{\mathbf{x_{T}}}$ ${\mathbf{x_{T}}}\rightarrow{\mathbf{x_{0}}}$ ${\mathbf{x_{0}}}\rightarrow{\mathbf{\hat{x}_{T}}}$ ${\mathbf{\hat{x}_{T}}}\rightarrow{\mathbf{x_{0}}}$
ADM-32 $98.9_{\pm 1.2}$ $50.4_{\pm 1.9}$ $16.4_{\pm 6.1}$ $33.4_{\pm 2.7}$
ADM-64 $100_{\pm 0.0}$ $59.0_{\pm 7.1}$ $8.6_{\pm 9.3}$ $57.5_{\pm 7.3}$
ADM-256 $100_{\pm 0.0}$ $23.2_{\pm 4.8}$ $4.1_{\pm 1.4}$ $23.9_{\pm 5.0}$
IF $99.8_{\pm 0.2}$ $52.6_{\pm 5.9}$ $3.1_{\pm 1.8}$ $50.0_{\pm 5.6}$
LDM $100_{\pm 0.0}$ $100_{\pm 0.0}$ $100_{\pm 0.0}$ $100_{\pm 0.0}$
DiT $100_{\pm 0.0}$ $90.7_{\pm 10.1}$ $99.5_{\pm 1.7}$ $90.7_{\pm 10.3}$
Table 8: Accuracy on assigning noise or inverted latent encoding to the generated image and vice versa. For pixel DMs, assigning starting noise to the generated images can be done by selecting the closest one with $l_{2}$ distance, while other assignments are impossible. For latent-space models (LDM and DiT), all the assignments can be done using the L2 distance. In Sec. E.2, we present that this property holds for different numbers of steps $T$.
5 On the effect of the first inversion steps
Building on our prior findings, which indicate that initial inversion steps predominantly contribute to the divergence between latent variables and Gaussian noise, we conclude the experiments with an evaluation of how we could mitigate those issues with access to ground truth inversion steps.
To this end, we generate $N=10000$ images with $T=50$ diffusion steps with DiT and IF models and, similarly as in Sec. 3.4, we substitute intermediate noise predictions $\{\mathcal{E}_{t}^{S}\}_{t=50\dots 1}$ during sampling.
First, in Tab. 9, we show that by replacing the approximations of the model’s predictions only for the first $5$ inversion steps (10%) with the ground truth ones, we obtain the latent variables with less than $50\%$, for DiT, and less than $15\%$, for IF, pixel correlation compared to standard latents. This reduction of less diverse areas in latents comes along with a decrease in the noise reconstruction error, calculated as $l_{2}$ distance $\frac{1}{N}\Sigma_{i}\lVert x_{T}^{[i]}-\hat{x}_{T}^{[i]}\rVert_{2}$, and Kullback–Leibler (KL) divergence to the Gaussian noise. However, it does not affect the image reconstruction.
Inversion DiT IF
predictions Pixel Reconstruction KL Div. Pixel Reconstruction KL Div.
replaced (%$\%T$) Corr. Image Noise $\mathcal{N}(0;\mathcal{I})$ Corr. Image Noise $\mathcal{N}(0;\mathcal{I})$
ø ($0\%$) $0.16$ $0.05$ $0.15$ $0.011$ $0.64$ $0.07$ $0.55$ $0.60$
$1$ ($2\%$) $0.16$ $0.05$ $0.15$ $0.011$ $0.61$ $0.07$ $0.51$ $0.56$
$1\dots 3$ ($6\%$) $0.10$ $0.05$ $0.08$ $0.004$ $0.15$ $0.07$ $0.19$ $0.06$
$1\dots 5$ ($10\%$) $0.07$ $0.05$ $0.06$ $0.002$ $0.09$ $0.07$ $0.12$ $0.03$
$1\dots 10$ ($20\%$) $0.04$ $0.05$ $0.03$ $0.001$ $0.05$ $0.07$ $0.07$ $0.01$
Noise ($100\%$) $0.04$ $0.0$ $0.0$ $0.05$ $0.0$ $0.0$
Table 9: Impact of DDIM Inversion approximation error in the first steps of the noising process on latent correlations, reconstruction errors, and the KL divergence between inversions and Gaussian noise. By starting the inversion procedure after $10\%$ of the image corruption process, we significantly remove image structures from latents and drastically reduce their divergence from noise.
Next, we evaluate how replacing the diffusion model’s predictions affects the diversity and text alignment of generated outputs. As shown in Fig. 7 for IF, swapping DM’s approximated predictions with denoising outputs during the first $3$ inversion steps reduces the difference in diversities for images resulting from latents and those generated from noise by $50\%$. This reduction is observed across both distance-based metrics (DreamSim, LPIPS) and similarity-based metrics (SSIM, DINO). Additionally, in Fig. 7, we show for both models that replacing the first $3$ DM predictions during inversion results in a significant decrease in modified generations alignment to the source prompt ($P_{S}$).
@@ 5 On the effect of the first inversion steps
Figure 7: Impact of first DDIM inversion errors on latent generations’ diversity and text-alignment. For IF model (right), replacing the first inversion steps (denoted as $r$) and denoising with a different prompt leads to more diverse generations against the source images $I_{S}$. For DiT (left), we show that replacing increases generations’ alignment to the target prompts, as indicated by the larger CLIP-T value for $P_{T}$ and the smaller one for $P_{S}$. We present details of the experiment in Sec. F.2.
In Sec. F.1, we show that replacing predictions in the first $3\%$ of steps also noticeably improves the quality of images generated from interpolations of final latents.
@@ 6 Conclusions
In this work, we analyze how errors introduced by the DDIM inversion algorithm influence the properties of latent image representations. Our findings reveal that the inverted latents exhibit systematic deviations from a Gaussian distribution, particularly in regions corresponding to smooth image surfaces. We trace this effect to the higher inversion error in these areas, which predominantly arise during the early noising steps. This discrepancy leads to practical limitations such as reduced diversity and lower text alignment to the conditioning prompts, or worse quality interpolation and editing in situations where latents are used as a starting point of the diffusion process. Our conclusions suggest that reducing the error difference for plain and non-plain image surfaces during the first inversion steps could mitigate these constraints, enhancing the usability of latents in generative applications and pointing additional direction for future research.
Supplementary Material
Appendix A On approximation error in DDIM Inversion
@@
Denoising Diffusion Probabilistic Models (DDPMs [14]) generate samples by reversing the forward diffusion process, modeled as a Markov Chain, where a clean image $x_{0}$ is progressively transformed to white Gaussian noise $x_{T}$ in $T$ diffusion steps. A partially noised image $x_{t}$, which serves as an intermediate object in this process, is expressed as
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$ (7)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, $\alpha_{t}=1-\beta_{t}$, $\epsilon_{t}\sim\mathcal{N}(0,\mathcal{I})$, and $\{\beta_{t}\}_{t=1}^{T}$ is a variance schedule, controlling how much of the noise is contained in the image at the specific step $t$.
@@
Song et al. [44] reformulate the diffusion process as a non-Markovian, which leads to a speed-up of the sampling process. While previously obtaining a less noisy image at $x_{t}$ required all past denoising steps from $T$ till $(t+1)$, this approach allows skipping some steps during sampling. More precisely, the backward diffusion process is defined as a combination of predictions of image $x_{0}$, next denoising step $x_{t}$, and random noise:
$x_{t-1}=$ $\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{\theta}^{(t)}(x_{t})}{\sqrt{\bar{\alpha}_{t}}}\right)$ (9)
$+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)}(x_{t})+\sigma_{t}z_{t}$
with $\sigma_{t}=\eta\sqrt{\beta_{t}(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}$, and $z_{t}\sim\mathcal{N}(0,\mathcal{I})$.
@@
Appendix B Most probable triangles
In Sec. 3.3, we analyze where the latent encodings are distributed in relation to the initial noise and generated samples. In this section, we describe the methodology for obtaining such triangles, as well as investigate how those relations change depending on a number of diffusion steps $T$.
B.1 Methodology

### v3

@@ 4.2 Diversity and quality of image edition
Text-to-image diffusion models with DDIM inversion are often used to modify images [14, 38, 12, 19]. However, knowing that DDIM latents are less diverse in plain areas, we hypothesize that using them as a starting point might reduce the diversity and quality of the edited samples. To evaluate this, we leverage DiT [42] and IF [52] as conditional DMs with $T=50$ diffusion steps. For each model, we construct two sets of $1280$ randomly selected prompts (ImageNet classes for DiT, COCO-30K [31] prompts for IF): (1) source prompts $P_{S}$, used during the generation and inversion, and (2) target prompts $P_{T}$, used for edition. Using source prompts $P_{S}$, we generate images $I_{S}$ from Gaussian noise ${\mathbf{x_{T}}}$ and invert them back into the latents ${\mathbf{\hat{x}_{T}}}$. Next, we regenerate images $\hat{I}_{T}$ from the latents, changing the conditioning to the target prompts $P_{T}$. We compare the edits with ground truth targets $I_{T}$ generated from the original noise ${\mathbf{x_{T}}}$ with $P_{T}$. In fig. 6, we present the drawback of leveraging latent variables as starting points for the denoising, where the $I_{S}$ images’ structures in the latents limit editing performance at those locations in $\hat{I}_{T}$ images.
Figure 6: Image edition by reversing the original image with a source prompt and reconstructing it with a target one. DDIM Inversion produces less diverse image changes when the manipulation is related to plain regions in source images, contrary to when using ground-truth Gaussian noise. Replacing the first steps of DDIM Inversion with forward diffusion leads to more editable latents.
@@
In table 10, we quantitatively measure this effect. First, we calculate the diversity of target generations ($I_{T}$,$\hat{I}_{T}$) against source images ($I_{S}$). We use DreamSim distance [11], LPIPS [61], SSIM [57], and cosine similarity of DINO features [5] to measure the distance between two sets of generations. The experiment shows that edits resulting from latent encodings $\hat{I}_{T}$ are characterized by higher similarity (SSIM, DINO) and lower diversity (DreamSim, LPIPS) relative to starting images $I_{S}$ than the one resulting from noises $I_{T}$. At the same time, in the bottom rows of table 10,
@@
Diversity DreamSim $\uparrow$ $0.71_{\pm 0.12}$ $0.68_{\pm 0.13}$ $0.67_{\pm 0.10}$ $0.61_{\pm 0.11}$
against $I_{S}$ LPIPS $\uparrow$ $0.59_{\pm 0.12}$ $0.57_{\pm 0.12}$ $0.38_{\pm 0.11}$ $0.33_{\pm 0.11}$
SSIM $\downarrow$ $0.23_{\pm 0.13}$ $0.26_{\pm 0.14}$ $0.34_{\pm 0.15}$ $0.41_{\pm 0.15}$
DINO $\downarrow$ $0.17$ $0.22$ $0.34$ $0.42$
Text CLIP-T ($P_{S}$) $\downarrow$ $0.465$ $0.480$ $0.273$ $0.649$
alignment CLIP-T ($P_{T}$) $\uparrow$ $0.689$ $0.568$ $0.353$ $0.614$
@@
Table 10: Diversity of editions (generations from noise ${\mathbf{x_{T}}}$ and latents ${\mathbf{\hat{x}_{T}}}$, conditioned on target prompt) in relation to source images $I_{S}$ and their alignment with source and target prompts. The arrows ($\uparrow$/$\downarrow$) indicate greater generation diversity and higher text alignment to the target prompt.
we show that the correlations occurring in the latent encodings induce lower performance in text-alignment to target prompts $P_{T}$, which we measure by calculating cosine similarity between text embeddings and resulting image embeddings, both obtained with the CLIP [44] encoder. For the DiT model, our experiment indicates that generations from latents are as similar to the target prompt from the regeneration as to the source one.
@@
Figure 7: Replacing first DDIM inversion steps with forward diffusion improves editions’ pixel diversity (a) and prompt-alignment (b). For IF and DiT models, diversity of generations can be improved by leveraging the forward diffusion process in the first inversion steps, and denoising resulting latents with a different prompt. Additionally, we observe a boost in generations’ alignment to the target prompts, as indicated by the larger CLIP-T value for $P_{T}$ and the smaller one for $P_{S}$. We present details of the experiment in appendix J.
We further evaluate how replacing the first inversion steps with the forward diffusion affects the diversity and text alignment of generated outputs. As shown in fig. 7a, swapping the first $4\%$ of DDIM inversion steps with forward diffusion improves the diversity of images generated from latents almost to the level of the samples from original noise. At the same time, as presented in fig. 7b, replacing the first inversion steps leads to a significant decrease in generations’ alignment to the source prompt ($P_{S}$) and an increase in similarity to the target prompt ($P_{T}$). This can also be observed in visualisations (fig. 6).
Conversely, replacing 4% of model predictions with a forward diffusion makes the process faster by around $3.9\%$. Although at the same time, intuitively, replacing the predicted noise with a random one should degrade reconstruction quality, recent studies [6, 33, 30, 10] have shown that the backward diffusion final steps do not contribute additional generative information, functioning instead as a data-agnostic denoising process. Therefore, substituting noise only in the initial steps of the inversion (e.g., the first 4%) allows us to enhance the quality of the latent representations while maintaining a reconstruction error at the level of the standard DDIM inversion. This means that the error between the ground truth prediction and DDIM approximation (eq. 4) is comparable to the error introduced with random Gaussian noise of the adjusted magnitude. Additionally, as presented in appendix K, our approach enables stochastic image editing, producing diverse manipulations of the same image.
On top of the experiments presented in this section, in the appendices, we provide additional studies comparing noise-to-image and latent-to-image mappings (appendix L) and the relation between noise and samples during training (appendix M).
@@ 5 Conclusions
In this work, we analyze how errors introduced by the DDIM inversion algorithm influence the properties of latent image representations. Our findings reveal that the inverted latents exhibit systematic deviations from a Gaussian distribution, particularly in regions corresponding to smooth image surfaces. We trace this effect to the higher inversion error in these areas, which predominantly arises during the early noising steps. This discrepancy leads to practical limitations such as reduced diversity and lower text alignment to the conditioning prompts, or worse quality interpolation and editing in situations where latents are used as a starting point of the diffusion process. To mitigate this issue, we introduce a simple yet effective solution, where we decorrelate latents by replacing the initial steps of the inversion procedure with forward diffusion. Our approach significantly improves the quality of samples in the practical use cases of DDIM inversion.
Appendix
[…]
[app]l1
Contents
In the Appendix we first outline the limitations (A) of our experiments, discuss the broader impact (B) of the introduced method, and list the computational resources (C) used. We then describe, in detail, the DDIM approximation error (D), which effects we study in this work and the consequences it has across varying numbers of sampling steps (E).
Next, we describe our experimental procedures: measuring the angles of the noise–image–latent triangle (F), identifying plain-regions pixels (G) in the image, and computing the inversion error (H). In section I, we provide the full pseudocode for the proposed algorithm.
We demonstrate how the technique enhances image diversity and prompt alignment (J) and how it enables stochastic editing of real images (K). Finally, we compare Gaussian noise and DDIM latents in their mappings to images (L) and track how these relationships evolve during diffusion training (M). In section N, we include additional qualitative results for the experiments, especially when introduced method might fail, and how it performs in the task of image interpolation.
Appendix A Limitations
In this work, we analyzed the relation between the random noise, image generations and their latent encodings obtained through DDIM Inversion. While our studies focused on DDIM approximation error from Song et al. [50], there exist other solvers and inversion methods, as described in section 2, bringing their own advantages and limitations. The error of DDIM inversion strongly depends on the number of steps with which it is performed. In particular, performing the process very granularly, e.g., using $T=1000$ steps, can result in strong suppression of the correlation. Nevertheless, the default number of steps we have chosen, i.e., 100, is, according to previous works [18, 12, 24], a practical choice as a good balance between the reconstruction error and the speed of the algorithm. In appendix E, we present that the latents exhibit correlations when using more sampling steps.
The observations from our analytical experiments (correlations in table 2, angles in table 3, variance in table 4, interpolations in fig. 5) generalize well to all tested diffusion models, but are less evident in the LDM model trained on the CelebA-HQ images. We attribute this exemption to the specificity of the dataset on which the model was trained - photos with centered human faces, usually with uniform backgrounds. We believe that, unlike models trained on a larger number of concepts, the process of generating faces with uniform backgrounds is more stable and introduces little detail in subsequent steps, making the difference in approximation error for plain and non-plain areas less significant.
As mentioned in section 2, the DDIM inversion error can be detrimental when using a small number of steps. Even though the solution proposed in this work (involving forward diffusion in first inversion steps) drastically removes correlations in latents and, thus, improves image interpolation and editing, it does not improve the numerical inversion error resulting from using small number of steps. In our experiments with 50 steps that are commonly used for edition, we show no significant drawbacks. However, in the extreme cases, using our method in even a single step, might result in the loss of information necessary for correct image reconstruction, hence it may be less preferred than standard DDIM inversion in some cases. In section N.1, we present failure cases for introduced solution.
Appendix B Broader impact
As our work is mostly analytical, we do not provide new technologies that might have a significant societal impact. However, our solution for improving the accuracy of DDIM inversion has potential implications that extend beyond technical advancements in diffusion models. As our approach enables more precise image editing it could be misused to advance image manipulation techniques. The enhanced interpolation quality could make synthetic content more convincing and harder to detect.
Appendix C Compute resources
For the experiments, we used a scientific cluster consisting of 110 nodes with CrayOS operating system. Each node is powered by 288 CPU cores, stemming from 4 NVIDIA Grace processors, each with 72 cores and a clock speed of 3.1 GHz. The nodes are equipped with substantial memory, featuring 480 GB of RAM per node. For GPU acceleration, each node in the cluster consists of 4 NVIDIA GH200 96GB GPUs with 120 GB of RAM and 72 CPUs per GPU.
Almost all of the experiments we perform are based on performing a sampling process using a diffusion model from noise, performing DDIM inversion and, possibly, image reconstruction from the latent, where each of these processes takes the same number of steps, hence the same number of GPU-hours on average. As our experiments differ in terms of number of sampling steps and images to generate, we provide average GPU time per one sampling step per batch (with $B$ denoting batch size) for each model as following: ADM-32 ($B=256$): 0.054s, ADM-64 ($B=128$): 0.093s, ADM-256 ($B=64$): 0.901s, LDM ($B=128$): 0.273s, DiT ($B=128$): 0.104s, IF ($B=64$): 0.609s. Note that some experiments, such as analyzing inversion approximation errors per step (fig. 4) or sampling from noise interpolations (fig. 5), involves performing the procedures several times.
Taking into consideration all the experiments described in the main text of this work, fully reproducing them takes roughly 95 GPU hours. However, considering the prototyping time, preliminary and failed experiments, as well as the fact that most of the experiments must be performed sequentially (e.g., inversion after image generation, image reconstruction after inversion), the overall execution time of the entire research project is multiple times longer.
Appendix D Approximation error in DDIM Inversion
@@
Denoising Diffusion Probabilistic Models (DDPMs [16]) generate samples by reversing the forward diffusion process, modeled as a Markov Chain, where a clean image $x_{0}$ is progressively transformed to white Gaussian noise $x_{T}$ in $T$ diffusion steps. A partially noised image $x_{t}$, which serves as an intermediate object in this process, is expressed as $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$ where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, $\alpha_{t}=1-\beta_{t}$, $\epsilon_{t}\sim\mathcal{N}(0,\mathcal{I})$, and $\{\beta_{t}\}_{t=1}^{T}$ is a variance schedule, controlling how much of the noise is contained in the image at the specific step $t$.
@@
Song et al. [50] reformulate the diffusion process as a non-Markovian, which leads to a speed-up of the sampling process. While previously obtaining a less noisy image at $x_{t}$ required all past denoising steps from $T$ till $(t+1)$, this approach allows skipping some steps during sampling. More precisely, the backward diffusion process is defined as a combination of predictions of image $x_{0}$, next denoising step $x_{t}$, and random noise (with $\sigma_{t}=\eta\sqrt{\beta_{t}(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}$ and $z_{t}\sim\mathcal{N}(0,\mathcal{I})$):
$x_{t-1}=$ $\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{\theta}^{(t)}(x_{t})}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)}(x_{t})+\sigma_{t}z_{t}.$ (7)
@@
Appendix E Correlations vs. inversion steps
In this work, for the inversion process, we leverage the DDIM sampler with either $T=50$ or $T=100$ sampling steps. This choice aligns with prior works [18, 12, 24] in image edition domain, where the authors used from $50$ up to $200$ inversion steps as a proper balance between reconstruction quality and short algorithm runtime.
However, as described in Hong et al. [18], the naïve DDIM inversion procedure [50] can be reinterpreted as solving the forward diffusion ordinary differential equation (ODE) in reverse order (along the time axis) with Euler method. With this reformulation, the inversion is correct under the assumption that, with $dt$ being step size, noise predictions in $t$ and $t+dt$ steps are almost exact, thus works only when performing with many iterations.
Since the presence of image structures in DDIM latents can depend on the number of steps with which the inversion is performed, in table 11 we show for ADM-64, DiT, and IF models that the observations we presented in this work generalize to cases where the number of inversion steps is several times greater (i.e., $T=1000$, as performed during the training). Additionally, in fig. 9 we show qualitatively by plotting the absolute error between starting Gaussian noise and DDIM latents, that also for a large number of steps, the uniform areas on the image contribute more significantly to overall inversion error.
Object Model
ADM-64 DiT IF
Noise ${\mathbf{x_{T}}}$ (baseline) $0.039_{\pm.00}$ $0.039_{\pm.00}$ $0.039_{\pm.00}$
DDIM Latent $\mathbf{\hat{x}_{T=10}}$ $0.416_{\pm.03}$ $0.297_{\pm.01}$ $0.783_{\pm.01}$
DDIM Latent $\mathbf{\hat{x}_{T=25}}$ $0.242_{\pm.02}$ $0.203_{\pm.02}$ $0.698_{\pm.02}$
DDIM Latent $\mathbf{\hat{x}_{T=50}}$ $0.177_{\pm.02}$ $0.144_{\pm.02}$ $0.608_{\pm.02}$
DDIM Latent $\mathbf{\hat{x}_{T=100}}$ $0.133_{\pm.01}$ $0.106_{\pm.02}$ $0.500_{\pm.02}$
DDIM Latent $\mathbf{\hat{x}_{T=250}}$ $0.108_{\pm.01}$ $0.078_{\pm.01}$ $0.366_{\pm.02}$
DDIM Latent $\mathbf{\hat{x}_{T=500}}$ $0.100_{\pm.01}$ $0.069_{\pm.01}$ $0.294_{\pm.02}$
DDIM Latent $\mathbf{\hat{x}_{T=1000}}$ $0.095_{\pm.01}$ $0.065_{\pm.01}$ $0.249_{\pm.02}$
Table 11: Latent encodings resulting from the DDIM Inversion exhibit correlations, even when the procedure is performed with a lot of steps. By dividing latent encodings into $8\times 8$ patches and calculating inside the mean of top-$20$ absolute Pearson correlation coefficients, we show that DDIM latents are correlated substantially higher than Gaussian noise, even when using $T=1000$ inversion steps.
Figure 9: Approximation errors in DDIM inversion are significantly higher for plain image surfaces than for the rest of the image. Even when using $T=1000$ steps, we observe image structures in resulting latents, which can be attributed to less diverse outputs for uniform image regions.
Appendix F Most probable triangles
In section 3.2, we analyze where the latent encodings are distributed in relation to the initial noise and generated samples. In this section, we describe the methodology for obtaining such triangles, as well as investigate how those relations change depending on a number of diffusion steps $T$.
F.1 Methodology